A particle with in SHM has an equation




X=0.75Sin65t cm. What is:




A. the frequency of the oscillation




B. The amplitude of the oscillation




C. The time period of the oscillation










1
Expert's answer
2022-01-16T13:12:39-0500

"x=0.75\sin65t"

"x = A\sin\omega t"

"A = 75 cm\text{ (amplitude)}"

"\omega = 65\text{ (cyclic frequency)}"

"A.f = \frac{\omega}{2\pi}= \frac{65}{2\pi}= 10.35 Hz"

"B. A = 75 cm = 0.75 m"

"C.T = \frac{1}{f}=\frac{1}{10.35}=0.097 s"


"\text{Answer:}"

"A.f = 10.35 Hz"

"B. A = 0.75 m"

"C.T = 0.097 s"